Given the following information:

Gov't private expenditure: 40

Gov't expenditure on goods and services: 17.50

Investments: 10.75

Export income from abroad: 12.50

Import income paid to abroad: 11.75

Taxes on expenditure: 9

Capital Consumption: 6.20

General subsidies: 0.75

1. Find GDP, NDP and National income

2. What method of computing is used?








1
Expert's answer
2021-06-28T17:26:04-0400

(a)

GDP = consumption + investment+ government spending+ (exports - imports)

"GDP = 6.20 + 10.75 + 40 + 0.75 + 17.50 + 9 + ( 12.5 - 11.75)"

"GDP = 84.95"

(b)

NDP = GDP- Depreciation

Depreciation"\implies" capital consumption

"NDP = 84.95 - 6.20"

"= 78.75"

(c)

National Income = consumption+ government spending + investment + net exports +Foreign production by resident of country- Domestic production by resident of other country.

"= 6.20 + 10.75+ 40 + 0.75+ 17.50 + 9 + (12.50. -11.75). +0-0"

"= 84.95"

2.

The method used in computation is the expenditure method. In this method we take the summation of total expenditure of private sector and government sector in an economy.